A researcher computes a one-way within-subjects ANOVA in which k = 3 and n = 10. What is the critical value for this test at a .05 level of significance?

a.

3.16

b.

4.41

c.

3.55

d.

5.48
Answer #1

Numerator degrees of freedom = k - 1 = 3 - 1 = 2

Denominator degrees of freedom = n - k = 10 - 3 = 7

Hence,

Critical value at 0.05 significance level = 4.41

Option B is correct.